OpenStudy (rainbow_rocks03):

Which best describes the spread of a set of data that has an interquartile range of 18 and a mean absolute deviation of 10? A. The average distance of all data values from the mean is 18 and the middle 50% of data values has a range of 10. B. The average distance of all data values from the mean is 14 and the middle 50% of data values has a range of 8. C. The average distance of all data values from the mean is 14 and the middle 50% of data values has a range of 28. D. The average distance of all data values from the mean is 10 and the middle 50% of data values has a range of 18.

OpenStudy (rainbow_rocks03):

@tkhunny

OpenStudy (tkhunny):

You still haven't done the other problem. Please don't post problems with no work shown. What say you?

OpenStudy (rainbow_rocks03):

what say you? what does that mean?

OpenStudy (rainbow_rocks03):

this is from the other problem. 38 - 19 = 19 27 - 19 = 8 21 - 19 = 2 19 - 16 = 3 19 - 10 = 9 19 - 2 = 17 19 + 8 + 2 + 3 + 9 + 17 = 58 58 / 6 ≈ 9.7

OpenStudy (tkhunny):

Why in the world would you post that here? So nice to be able to see your work. Very good. Keep that up.

OpenStudy (rainbow_rocks03):

now can you help me on THIS problem and TY

OpenStudy (tkhunny):

Mean Absolute Deviation = Mean of the Distance from the Mean. It is calculated as 10. Shall we pick 18 or 14? I think not.

OpenStudy (rainbow_rocks03):

you think not what?

OpenStudy (rainbow_rocks03):

@tkhunny

OpenStudy (tkhunny):

It's a complete sentence, given the context. You have to do some thinking. Go!